OpenStudy (anonymous):

Determine whether the following infinite series are convergent or divergent. a. 1 + ½ + 1/6 + 1/24 +…..

OpenStudy (zarkon):

use the ratio test...(also this is a well known series)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

with the ratio test i got 1, is this correct?

OpenStudy (zarkon):

no

OpenStudy (zarkon):

it looks like your series is \[\sum_{n=1}^{\infty}\frac{1}{n!}\]

OpenStudy (zarkon):

the ratio test will give a value of 0 notice that \(\displaystyle\sum_{n=1}^{\infty}\frac{1}{n!}=e-1\)